# What is a Schiff moment anyway?

### v2

@@ Abstract
Schiff moments of atomic nuclei are of considerable interest to experiments searching for undiscovered new physics that breaks time-reversal symmetry. I develop a simple picture of the Schiff moment of a charge distribution, and discuss the interaction of the Schiff moment of a nucleus with the field produced by an electron in an atom.
@@
The purpose of this note is two-fold: first, to provide a simple yet accurate description of the Schiff moment of a charge distribution, and second, to illustrate the interesting physics of the interaction of a nuclear Schiff moment with electrons in atoms. This note may be useful to undergraduates or graduate students interested in precision searches for physics beyond the Standard Model of particle physics. Perhaps it may be useful as well to students and instructors of electromagnetism courses, as an interesting example of basic electrostatics that can be connected to contemporary research.
@@ I.   Electron interaction with a nuclear charge distribution
Here $k=-e/4\pi\epsilon_{0}$ for ease of notation. The $E0,E1,E2$ and $E3$ interactions are traditionally called the electric monopole, dipole, quadrupole and octupole interactions respectively. Each of these is a product of a nuclear multipole moment (in round brackets) that depends on $\vec{R}$, and a field produced by the electron (in square brackets) that depends on $\vec{r}$. The gradients of $1/r$ appearing above are derived in the Appendix. The symmetrized 3rd rank tensor formed from the 2nd rank tensor $A_{ij}$ and the vector $B_{\ell}$ is compactly written as $A_{\{ij}B_{\ell\}}=A_{ij}B_{\ell}+A_{\ell i}B_{j}+A_{j\ell}B_{i}$.
In the $E2$ and $E3$ interactions, there are two kinds of fields produced by the electron. The so-called “contact” terms contain $\delta(\vec{r})$ or its derivatives [15, 16], whereas the non-contact terms are traceless tensors. Let us focus on the contact $E3$ term
@@
The quantity $\vec{\mathscr{S}}_{0}=\frac{1}{10}\int d^{3}R\,\rho_{n}(\vec{R})\,R^{2}\,\vec{R}$ is the first piece of the Schiff moment of the charge distribution $\rho_{n}(\vec{R})$. I have also defined $\vec{\mathcal{F}}=\frac{e}{\epsilon_{0}}\vec{\nabla}\delta(\vec{r})$, the field from the electron that interacts with the nuclear Schiff moment – this will be discussed in more detail in Section IV.
@@ II.   Shielding of the nuclear electric dipole moment in an atom
  4. (iv)
    In a quantum mechanical system with angular momentum, such as a nucleus with spin $\vec{I}$, the Schiff moment is parallel to the total angular momentum, $\vec{\mathscr{S}}\propto\vec{I}$. This property follows from the Wigner-Eckart theorem [18].
@@ III.   Charge distribution with a Schiff moment
For concreteness, here is a simple example of a charge distribution with a nonzero Schiff moment. Consider a charge distribution confined to the surface of a sphere of radius $b$. Let the charge per unit area on the surface be $\sigma(\theta,\phi)=\frac{e}{4\pi b^{2}}\left(1+\eta\cos\theta\right)$, where $\theta,\phi$ are the usual spherical polar angles. Let us assume the dimensionless constant $\eta\ll 1$. The moments of this distribution are
@@
Since the charge distribution has azimuthal symmetry, the $X$ and $Y$ components of the EDM and Schiff moment are zero, and only the $Z$ components are nonzero. This distribution has both a monopole and a dipole moment, and so the centre of charge is not at the centre of the sphere, but at $\Delta\vec{R}=(0,0,\frac{\eta b}{3})$ instead. The $Z$ component of $\vec{\mathscr{S}}$ is easily calculated using Eq. (9) to be $\mathscr{S}_{Z}=\frac{1}{30}\eta eb^{3}-\frac{1}{18}\eta eb^{3}=-\frac{1}{45}\eta eb^{3}$.
IV.   The $\mathcal{F}$-field
Having seen how the definition of the Schiff moment arises naturally from the electron-nucleus electrostatic interaction, let us now consider the electron-generated field that couples to the nuclear Schiff moment. The definition of $\vec{\mathcal{F}}$ in Equation (8), which involves the gradient of a delta-function, is a rather singular mathematical object. To gain physical understanding without getting lost in mathematical subtleties, let us replace the $\delta(\vec{r})$ in the definition of $\vec{\mathcal{F}}$ with its smooth regularized version described in the Appendix, with a regularization length scale $a$. The gradient of the delta-function is then
@@ IV.   The $\mathcal{F}$-field
where the dimensionless function $g(r)=\frac{15}{4\pi}\left(1+\frac{r^{2}}{a^{2}}\right)^{-7/2}$ is defined for convenience. Thus a finite approximation of the $\mathcal{F}$-field produced by the electron is $\vec{\mathcal{F}}=-\frac{e\vec{r}\,g(r)}{\epsilon_{0}a^{5}}$. Fig. 1 shows an illustration of the $\mathcal{F}$-field from an electron in a classical orbit around a nucleus.
@@
Figure 1: The $\mathcal{F}$-field produced by an electron in an atom. This field is localized within a sphere of radius $\sim a$ (which we take to be much smaller than the size of an electron orbit), and is directed opposite to the electron’s position vector $\vec{r}$. The $\mathcal{F}$-field couples to the Schiff moment of the nucleus as described by Equation (8).
@@
But electrons do not move classically inside atoms, so it is instructive to calculate the $\mathcal{F}$-field produced by an electron in a hydrogenic atom. Consider an electron in a superposition of $ns$ and $n^{\prime}p_{z}$ states, where $n,n^{\prime}$ denote hydrogenic principal quantum numbers: ${\left|{\psi}\right\rangle}=c_{s}{\left|{ns}\right\rangle}+c_{p}{\left|{n^{\prime}p_{z}}\right\rangle}$, with complex coefficients $c_{s}$ and $c_{p}$. The spatial wavefunctions of the atomic states are ${\left\langle{\vec{r}}|{ns}\right\rangle}=\mathfrak{R}_{ns}(r)Y_{00}(\theta\phi)$ and ${\left\langle{\vec{r}}|{n^{\prime}p_{z}}\right\rangle}=\mathfrak{R}_{n^{\prime}p}(r)Y_{10}(\theta\phi)$, where $\mathfrak{R}_{n\ell}(r)$ are hydrogenic radial functions and $Y_{\ell m}(\theta\phi)$ are spherical harmonics. Note that ${\left\langle{ns}\right|}\mathcal{F}_{z}{\left|{ns}\right\rangle}$ and ${\left\langle{n^{\prime}p}\right|}\mathcal{F}_{z}{\left|{n^{\prime}p}\right\rangle}$ both vanish, since $\mathcal{F}_{z}$ is an odd-parity operator sandwiched between definite-parity wavefunctions, which makes the expectation value integral equal zero. Therefore only the cross terms between $\left|{ns}\right\rangle$} and $\left|{n^{\prime}p}\right\rangle$} contribute, and so ${\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}=2\,\mathrm{Re}(c_{p}^{*}c_{s}){\left\langle{n^{\prime}p}\right|}\mathcal{F}_{z}{\left|{ns}\right\rangle}$. Defining $\xi=\mathrm{Re}(c_{p}^{*}c_{s})$, the expectation value of the $z$-component of $\vec{\mathcal{F}}$ in the electron state $\left|{\psi}\right\rangle$} is
@@
$\begin{split}{\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}&=2\xi{\left\langle{n^{\prime}p}\right|}\mathcal{F}_{z}{\left|{ns}\right\rangle}\\ &=-\frac{2\xi e}{\epsilon_{0}a^{5}}\int_{0}^{\infty}4\pi r^{2}dr\,\mathfrak{R}_{ns}(r)\mathfrak{R}_{n^{\prime}p}(r)\,rg(r)\int_{0}^{2\pi}d\phi\int_{0}^{\pi}d\theta\,\sin\theta\,Y_{00}(\theta\phi)\cos\theta\,Y_{10}(\theta\phi)\\ &=-\frac{2\xi e}{\sqrt{3}\epsilon_{0}a^{5}}\int_{0}^{r_{1}}4\pi r^{2}dr\,\mathfrak{R}_{ns}(r)\mathfrak{R}_{n^{\prime}p}(r)\,rg(r).\end{split}$ (12)
@@
Now imagine that we set the length scale $a$ to be much smaller than the typical size of an atom, so that inside a sphere of radius $a$ we can approximate $\mathfrak{R}_{ns}(r)\approx\alpha$ and $\mathfrak{R}_{n^{\prime}p}(r)=\beta r$, their values close to the origin. The constants $\alpha$ and $\beta$ can be obtained from the leading-order series approximation of hydrogenic wavefunctions, and their values are $\alpha=2\sqrt{\frac{Z^{3}}{n^{3}a_{0}^{3}}}$ and $\beta=\frac{2Z}{3a_{0}}\sqrt{\frac{Z^{3}}{n^{\prime 3}a_{0}^{3}}\left(1-\frac{1}{n^{\prime 2}}\right)}$, where $Z$ is the nuclear charge and $a_{0}$ is the Bohr radius.
Thus the $\mathcal{F}$-field produced by the electron is
@@
Remarkably, the integral $\int_{0}^{r_{1}}4\pi r^{2}dr\,r^{2}g(r)=3a^{5}$ whenever $r_{1}\gtrsim a$, since the step-like function $g(r)$ only has appreciable support over a region of radius $\sim a$ around the origin. Hence ${\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}$ has a finite and well-defined value, regardless of our choice of specific values for the regularization length scale $a$ and the integral limit $r_{1}$ as long as $a\ll r_{1}\ll a_{0}/Z$. Thus
@@
If a nucleus with Schiff moment $\mathscr{S}_{z}\hat{z}$ interacts with electrons in the state $\left|{\psi}\right\rangle$}, then ${\left\langle{\psi}\right|}V_{cE3}^{\prime}{\left|{\psi}\right\rangle}=-\mathscr{S}_{z}{\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}$. The Schiff moment can be written in terms of the nuclear spin vector $\vec{I}$, as $\mathscr{S}_{z}=\frac{|\vec{\mathscr{S}}|}{|\vec{I}|}I_{z}$, due to the Wigner-Eckart theorem [18]. Thus, when the electron is in state $\left|{\psi}\right\rangle$}, the effective Hamiltonian for the nuclear spin degree of freedom is $H_{\mathrm{NSM}}=-I_{z}\frac{|\vec{\mathscr{S}}|}{|\vec{I}|}{\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}$. When the nuclear spin is in the eigenstate $\left|{I,m_{I}}\right\rangle$}, the Hamiltonian $H_{\mathrm{NSM}}\propto I_{z}$ leads to characteristic energy shifts ${\left\langle{I,m_{I}}\right|}H_{\mathrm{NSM}}{\left|{I,m_{I}}\right\rangle}\propto m_{I}\,{\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}$, which can be measured in experiments (see, e.g., [3, 4]).
Note some interesting properties of $\vec{\mathcal{F}}$. Its expectation value goes to zero if either $c_{s}$ or $c_{p}$ is zero, so only an electron in a superposition of $s$ and $p$ states generates a nonzero $\mathcal{F}$-field in the vicinity of the nucleus. The magnitude of $\vec{\mathcal{F}}$ also scales strongly with the nuclear charge $Z$. These are the reasons why current experiments searching for nuclear Schiff moments use heavy nuclei contained within electrically-polarized atoms or molecules (e.g., [3, 4, 5]), wherein the electrons are in superpositions of atomic $s$ and $p$ states around a nucleus with large $Z$.
In summary, a Schiff moment is a property of a charge distribution that combines aspects of its monopole, dipole, quadrupole and octupole moments. The definition of a nuclear Schiff moment, and the physics of its interaction with electrons in atoms, follow from elementary electrostatics.

### v3

@@ Abstract
Schiff moments of atomic nuclei are of considerable interest to experiments searching for undiscovered new physics that breaks time-reversal symmetry. I develop a simple picture of the Schiff moment of a charge distribution, and discuss the interaction of the Schiff moment of a nucleus with the electron in an atom.
@@
The purpose of this note is two-fold: first, to provide a simple yet accurate description of the Schiff moment of a charge distribution, and second, to illustrate the physics of the interaction of a nuclear Schiff moment with electrons in atoms. This note may be useful to undergraduates or graduate students interested in precision searches for physics beyond the Standard Model of particle physics. Perhaps it may be useful as well to students and instructors of electromagnetism courses, as an interesting example of basic electrostatics that can be connected to contemporary research. Readers interested in an alternative pedagogical introduction to the Schiff moment are referred to Section 1.2 of Ref. [15].
@@ I.   Electron interaction with a nuclear charge distribution
Here $k=-e/4\pi\epsilon_{0}$ for ease of notation. The $E0,E1,E2$ and $E3$ interactions are traditionally called the electric monopole, dipole, quadrupole and octupole interactions respectively. Each of these is a product of a nuclear multipole moment (in round brackets) that depends on $\vec{R}$, and a field generated by the electron (in square brackets) that depends on $\vec{r}$. The gradients of $1/r$ appearing above are derived in the Appendix. The symmetrized 3rd rank tensor formed from the 2nd rank tensor $A_{ij}$ and the vector $B_{\ell}$ is compactly written as $A_{\{ij}B_{\ell\}}=A_{ij}B_{\ell}+A_{\ell i}B_{j}+A_{j\ell}B_{i}$.
Within the $E2$ and $E3$ interaction potentials, there are two types of electronic terms. The so-called “contact” terms contain $\delta(\vec{r})$ or its derivatives [16, 17], whereas the non-contact terms involve traceless tensors formed from components of $\vec{r}$. Let us focus our attention on the contact $E3$ term
@@
The quantity $\vec{\mathscr{S}}_{0}=\frac{1}{10}\int d^{3}R\,\rho_{n}(\vec{R})\,R^{2}\,\vec{R}$ is the first piece of the Schiff moment of the charge distribution $\rho_{n}(\vec{R})$. I have also defined $\vec{\mathcal{F}}=\frac{e}{\epsilon_{0}}\vec{\nabla}\delta(\vec{r})$, the vector that couples to the nuclear Schiff moment – this will be discussed in more detail in Section IV.
@@ II.   Shielding of the nuclear electric dipole moment in an atom
In quantum mechanics, the Schiff moment $\vec{\mathscr{S}}$ is a vector operator in the space of nuclear states. In a nucleus with spin $\vec{I}$, the expectation value of the Schiff moment vector in any nuclear state $\left|{\Psi}\right\rangle$} is parallel to the expectation value of the angular momentum vector, $\left\langle{\vec{\mathscr{S}}}\right\rangle={\left\langle{\Psi}\right|}\vec{\mathscr{S}}{\left|{\Psi}\right\rangle}\propto{\left\langle{\Psi}\right|}\vec{I}{\left|{\Psi}\right\rangle}$. This property follows from the Wigner-Eckart theorem [19]. Note, however, that $\vec{I}$ is an operator that is odd under time-reversal (T) symmetry: a spinning nucleus with angular momentum $\vec{I}$ can be thought of as a nucleus with angular momentum $\vec{I}$ whose dynamics is time-reversed. Therefore, if $\left\langle{\vec{\mathscr{S}}}\right\rangle\neq 0$ in an angular momentum eigenstate of a nucleus, then the nuclear charge distribution is odd under T-symmetry. Which means that the nucleus spinning in one sense has a different charge distribution than the nucleus spinning in the opposite sense! This kind of unusual charge distribution can only arise (as far as we know) from new particles or forces that break T-symmetry, which is why a nuclear Schiff moment is useful as an indicator of physics beyond the Standard Model.
@@ III.   Charge distribution with a Schiff moment
For concreteness, here is a simple example of a charge distribution with a nonzero Schiff moment. Consider a charge distribution confined to the surface of a sphere of radius $b$. Let the charge per unit area on the surface be $\sigma(\theta,\phi)=\frac{e}{4\pi b^{2}}\left(1+\eta\cos\theta\right)$, where $\theta,\phi$ are the usual spherical polar angles. The surface charge density is illustrated in Figure 1. Let us assume the dimensionless constant $\eta\ll 1$. The moments of this distribution are
@@
Figure 1: An example of a charge distribution with a nonzero Schiff moment. The color bar shows the surface charge density $\sigma(\theta,\phi)\propto(1+\eta\cos\theta)$ on the sphere, with $\eta=0.1$. The direction of the Schiff moment of this charge distribution is indicated by the arrow.
Since the charge distribution has azimuthal symmetry, the $X$ and $Y$ components of the EDM and Schiff moment are zero, and only the $Z$ components of these vectors are nonzero. This distribution has both a monopole and a dipole moment, and so the centre of charge is not at the centre of the sphere, but at $\Delta\vec{R}=(0,0,\frac{\eta b}{3})$ instead. The $Z$ component of $\vec{\mathscr{S}}$ is easily calculated using Eq. (9) to be $\mathscr{S}_{Z}=\frac{1}{30}\eta eb^{3}-\frac{1}{18}\eta eb^{3}=-\frac{1}{45}\eta eb^{3}$.
IV.   The $\vec{\mathcal{F}}$ vector
Having seen how the definition of the Schiff moment arises naturally from the electron-nucleus electrostatic interaction, let us explore the physics of the electronic quantity $\vec{\mathcal{F}}$ that couples to the nuclear Schiff moment. The definition of $\vec{\mathcal{F}}$ in Equation (8), which involves the gradient of a delta-function, is a rather singular mathematical object. To gain physical understanding without getting lost in mathematical subtleties, let us replace the $\delta(\vec{r})$ in the definition of $\vec{\mathcal{F}}$ with its smooth regularized version described in the Appendix, using a regularization length scale $a$. The gradient of the delta-function is then
@@ IV.   The $\vec{\mathcal{F}}$ vector
where the dimensionless function $g(r)=\frac{15}{4\pi}\left(1+\frac{r^{2}}{a^{2}}\right)^{-7/2}$ is defined for convenience. Thus a non-singular approximation of $\vec{\mathcal{F}}$ generated by the electron is $\vec{\mathcal{F}}=-\frac{e\vec{r}\,g(r)}{\epsilon_{0}a^{5}}$. Fig. 1 shows an illustration of the $\vec{\mathcal{F}}$ vector field produced by an electron in a classical orbit around a nucleus.
@@
Figure 2: Illustration of the vector field $\vec{\mathcal{F}}$ produced by an electron in an atom. This vector field is localized and approximately constant within a sphere of radius $\sim a$ (which we can take to be much smaller than the size of an electron orbit), and is directed opposite to the electron’s position vector $\vec{r}$. The vector $\vec{\mathcal{F}}$ couples to the Schiff moment of the nucleus, as described by Equation (8).
But electrons do not move classically inside atoms. It is instructive to calculate the quantum mechanical expectation value of the $\vec{\mathcal{F}}$ operator for an electron in a simple hydrogenic atom. Without loss of generality, we can choose a coordinate system that is oriented along $\vec{\mathcal{F}}$, so let us simply calculate the expectation value of the $z$ component of $\vec{\mathcal{F}}$. First consider an electron in an $s$ state (orbital angular momentum $\ell=0$) with principal quantum number $n$. It turns out that ${\left\langle{ns}\right|}\mathcal{F}_{z}{\left|{ns}\right\rangle}=0$. This is a consequence of two facts: i) the operator $\mathcal{F}_{z}$ is a component of a polar vector, and ii) angular momentum eigenstates like have definite parity. Denoting the unitary parity-reversal operator by $P$, these facts are expressed by the relations $P\mathcal{F}_{z}P^{\dagger}=-\mathcal{F}_{z}$ and $P{\left|{ns}\right\rangle}=+{\left|{ns}\right\rangle}$. And so, making use of the unitary property $P^{\dagger}P=\mathbb{1}$,
${\left\langle{ns}\right|}\mathcal{F}_{z}{\left|{ns}\right\rangle}={\left\langle{ns}\right|}P^{\dagger}P\,\mathcal{F}_{z}\,P^{\dagger}P{\left|{ns}\right\rangle}=({\left\langle{ns}\right|}P^{\dagger})\,(-\vec{\mathcal{F}}_{z})\left(P{\left|{ns}\right\rangle}\right)=-{\left\langle{ns}\right|}\mathcal{F}_{z}{\left|{ns}\right\rangle}=0.$ (12)
Consider an electron in the state $\left|{n^{\prime}p_{z}}\right\rangle$} instead, where $n^{\prime}$ is the principal quantum number, and $p_{z}$ denotes the orbital with angular momentum $\ell=1$ and azimuthal quantum number $m_{\ell}=0$. Here too, since $P{\left|{p_{z}}\right\rangle}=-{\left|{p_{z}}\right\rangle}$, it is easy to verify that ${\left\langle{n^{\prime}p_{z}}\right|}\mathcal{F}_{z}{\left|{n^{\prime}p_{z}}\right\rangle}=0$, following the same lines as Equation (12). So far, it seems like neither $s$ states nor $p$ states generate any interesting $\vec{\mathcal{F}}$ for a nuclear Schiff moment to interact with.
@@
Now consider an electron in a superposition of $ns$ and $n^{\prime}p_{z}$ states, ${\left|{\psi}\right\rangle}=c_{s}{\left|{ns}\right\rangle}+c_{p}{\left|{n^{\prime}p_{z}}\right\rangle}$, with complex coefficients $c_{s}$ and $c_{p}$, and let us evaluate ${\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}$. Since ${\left\langle{ns}\right|}\mathcal{F}_{z}{\left|{ns}\right\rangle}$ and ${\left\langle{n^{\prime}p_{z}}\right|}\mathcal{F}_{z}{\left|{n^{\prime}p_{z}}\right\rangle}$ are both equal to zero, only the cross terms between $\left\langle{ns}\right|$} and $\left|{n^{\prime}p_{z}}\right\rangle$} contribute to the expectation value: ${\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}=2\,\mathrm{Re}(c_{p}^{*}c_{s}){\left\langle{n^{\prime}p_{z}}\right|}\mathcal{F}_{z}{\left|{ns}\right\rangle}$. Let us define $\xi=\mathrm{Re}(c_{p}^{*}c_{s})$ for compactness of notation. To evaluate the matrix element ${\left\langle{n^{\prime}p_{z}}\right|}\mathcal{F}_{z}{\left|{ns}\right\rangle}$, we need the spatial wavefunctions of the atomic states, which are ${\left\langle{\vec{r}}|{ns}\right\rangle}=\mathfrak{R}_{ns}(r)Y_{00}(\theta\phi)$ and ${\left\langle{\vec{r}}|{n^{\prime}p_{z}}\right\rangle}=\mathfrak{R}_{n^{\prime}p}(r)Y_{10}(\theta\phi)$, where $\mathfrak{R}_{n\ell}(r)$ are hydrogenic radial functions and $Y_{\ell m}(\theta\phi)$ are spherical harmonics. Therefore
@@
$\begin{split}{\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}&=2\xi{\left\langle{n^{\prime}p_{z}}\right|}\mathcal{F}_{z}{\left|{ns}\right\rangle}\\ &=-\frac{2\xi e}{\epsilon_{0}a^{5}}\int_{0}^{\infty}4\pi r^{2}dr\,\mathfrak{R}_{ns}(r)\mathfrak{R}_{n^{\prime}p}(r)\,rg(r)\int_{0}^{2\pi}d\phi\int_{0}^{\pi}d\theta\,\sin\theta\,Y_{00}(\theta\phi)\cos\theta\,Y_{10}(\theta\phi)\\ &=-\frac{2\xi e}{\sqrt{3}\epsilon_{0}a^{5}}\int_{0}^{r_{1}}4\pi r^{2}dr\,\mathfrak{R}_{ns}(r)\mathfrak{R}_{n^{\prime}p}(r)\,rg(r).\end{split}$ (13)
@@
Let us set the regularization length scale $a$ (which, recall is just a mathematical artifice) to be much smaller than the typical size of an electron orbital in an atom. For a hydrogenic atom with nuclear charge $Z$, the size of atomic orbitals is $\sim a_{0}/Z$, where $a_{0}$ is the Bohr radius. Setting $a\ll a_{0}/Z$, inside a sphere of radius $a$ we can approximate the wavefunctions by their leading-order Taylor expansion near the origin, $\mathfrak{R}_{ns}(r)\approx\alpha$ and $\mathfrak{R}_{n^{\prime}p}(r)=\beta r$. The constants $\alpha$ and $\beta$ are easily obtained from tables of hydrogenic wavefunctions, and their values are $\alpha=2\sqrt{\frac{Z^{3}}{n^{3}a_{0}^{3}}}$ and $\beta=\frac{2Z}{3a_{0}}\sqrt{\frac{Z^{3}}{n^{\prime 3}a_{0}^{3}}\left(1-\frac{1}{n^{\prime 2}}\right)}$.
Thus the $z$-component of the $\vec{\mathcal{F}}$ vector field produced by an electron in state $\left|{\psi}\right\rangle$} is
@@
Remarkably, the integral $\int_{0}^{r_{1}}4\pi r^{2}dr\,r^{2}g(r)=3a^{5}$ whenever $r_{1}\gtrsim a$, regardless of the exact value of $r_{1}$, since the step-like function $g(r)$ only has appreciable support over a region of radius $\sim a$ around the origin. Hence ${\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}$ has a finite and well-defined value, regardless of our choice of specific values for the regularization length scale $a$ and the integral limit $r_{1}$, so long as $a\ll r_{1}\ll a_{0}/Z$. All of which leads to the expression
@@
Note that the expectation value of $\mathcal{F}_{z}$ goes to zero if either $c_{s}$ or $c_{p}$ is zero, so only an electron in a superposition of $s$ and $p$ states generates nonzero $\vec{\mathcal{F}}$ in the vicinity of the nucleus. Therefore, electrically-polarized atoms or polar molecules, wherein electron eigenfunctions consist of linear combinations of $s$ and $p$ orbitals, are necessary in order to generate an $\vec{\mathcal{F}}$ vector field to couple to a nuclear Schiff moment. Also note that ${\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}\propto Z^{4}$ is a steeply rising function of $Z$. These are the reasons why current experiments searching for nuclear Schiff moments use high-$Z$ nuclei contained within electrically-polarized atoms or molecules (e.g., [3, 4, 5]).
When a nucleus in state $\left|{\Psi}\right\rangle$} interacts with an electron in state $\left|{\psi}\right\rangle$}, then the energy shift of the combined atomic state ${\left|{\Psi}\right\rangle}{\left|{\psi}\right\rangle}$ due to the Schiff moment interaction is
$\Delta E={\left\langle{\psi}\right|}{\left\langle{\Psi}\right|}\,V_{cE3}^{\prime}\,{\left|{\Psi}\right\rangle}{\left|{\psi}\right\rangle}=-{\left\langle{\Psi}\right|}\mathscr{S}_{z}{\left|{\Psi}\right\rangle}\,{\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}$ (16)
to first order in perturbation theory. If the nucleus is prepared in the spin eigenstate ${\left|{\Psi}\right\rangle}={\left|{I,m_{I}}\right\rangle}$, the expectation value of the nuclear Schiff moment operator is
${\left\langle{\Psi}\right|}\mathscr{S}_{z}{\left|{\Psi}\right\rangle}={\left\langle{I,m_{I}}\right|}\mathscr{S}_{z}{\left|{I,m_{I}}\right\rangle}=\frac{|\vec{\mathscr{S}}|}{|\vec{I}|}{\left\langle{I,m_{I}}\right|}I_{z}{\left|{I,m_{I}}\right\rangle}=\frac{|\vec{\mathscr{S}}|}{|\vec{I}|}m_{I},$ (17)
due to the Wigner-Eckart theorem [19]. The resulting characteristic energy shift of the atom, $\Delta E\propto m_{I}\,{\left\langle{\psi}\right|}\mathcal{F}_{z}{\left|{\psi}\right\rangle}$, is measurable in experiments (see, e.g., [3, 4]).
In summary, a Schiff moment is a property of a charge distribution that combines aspects of its electric monopole, dipole, quadrupole and octupole moments. The definition of a nuclear Schiff moment, and the physics of its interaction with electrons in atoms, follow from elementary electrostatics and quantum mechanics.